# Cacophony: An Improved Contrastive Audio-Text Model

### v2

@@ Abstract
Despite recent advancements in audio-text modeling, audio-text contrastive models still lag behind their image-text counterparts in scale and performance. We propose a method to improve both the scale and the training of audio-text contrastive models. Specifically, we craft a large-scale audio-text dataset containing 13,000 hours of text-labeled audio, using pretrained language models to process noisy text descriptions and automatic captioning to obtain text descriptions for unlabeled audio samples. We first train on audio-only data with a masked autoencoder (MAE) objective, which allows us to benefit from the scalability of unlabeled audio datasets. We then, initializing our audio encoder from the MAE model, train a contrastive model with an auxiliary captioning objective. Our final model, which we name Cacophony, achieves state-of-the-art performance on audio-text retrieval tasks, and exhibits competitive results on the HEAR benchmark and other downstream tasks such as zero-shot classification.
@@ I Introduction
Machine audition [1] involves developing algorithms and systems for machines to analyze and understand sound, covering tasks such as audio tagging, acoustic scene classification, music classification, and sound event detection. In recent years, there has been a general shift away from individual audio pattern recognition tasks and towards general-purpose audio representations pretrained on large-scale audio datasets. Pretrained Audio Neural Networks (PANNs)  [2] have played a significant role in this shift by demonstrating their versatility across various tasks and outperforming many advanced systems via fine-tuning.
@@
Modern approaches aim for robust performance in general-purpose audio understanding tasks without the need for task-specific fine-tuning, which offers more flexibility. These methods approach general audio understanding by linking text and audio modality, referred to as audio-text models. One approach to linking text and audio is through generating response text given a combination of an audio prompt and a text prompt [3, 4, 5]. For instance, Pengi [5] converts audio classification, retrieval, captioning, and audio question answering into a text generation task using audio and task-specific text prompts. Similarly, Qwen-Audio [6] addresses a variety of audio tasks through text generation but distinguishes itself by using text prompts consisting of hierarchical tag sequences inspired by Whisper [7]. Both Pengi and Qwen-Audio support multiple close-ended and open-ended audio tasks without the need for additional fine-tuning or task-specific extensions of the architecture. Another method of linking text and audio is contrastive learning. Pretrained contrastive models can also be applied directly to various downstream tasks without fine-tuning. For instance, contrastive models can be used for retrieval and classification by assigning a score that identifies the most probable text (or class label) from a predefined set of choices for a given audio input. Moreover, the learned text-audio representations from contrastive models offers the flexibility to use one modality during training and the other during inference, which can be applied in text-to-audio generation [8, 9] and language-guided source separation [10, 11].
@@
In this paper, we concentrate on improving contrastive audio-text models for general sounds through dataset scaling and modeling techniques. In contrast to recent advancements in image-text contrastive models, audio-text contrastive models tend to lag behind in dataset scale. For example, in the image domain, CLIP [12] utilizes 400 million image-text pairs, and the more recent SigLip [13] employs 3.6 billion image-text pairs. On the other hand, the largest publicly available text-audio datasets contain less than 100 thousand pairs [14, 15]. To address the issue of data scarcity, previous works have focused on collecting data from various sources in the wild, relying on natural language processing techniques to clean or filter out noisy captions. Wu et al. [16] curate the LAION-Audio dataset with 630K audio-text pairs together with 2 million keyword-to-text (K2C) recaptioned Audioset [17] to train LAION-CLAP. The K2C method uses a pretrained language model to generate captions from tags. However, the captions produced via K2C are restricted to the objects defined by the tags, offering limited descriptive details. K2C also risks making incorrect assumptions or introducing biases highlighted in [16]. Mei et al. [18] propose a multi-stage data filtering pipeline and utilize ChatGPT for cleaning text descriptions. Their Contrastive Language-Audio Pretraining (CLAP) models, trained on their WavCaps dataset, demonstrate superior performance in audio-text retrieval tasks compared to LAION-CLAP, despite a smaller-scale dataset. Such a text filtering pipeline reduces the amount of audio seen by the model, which could potentially result in reduced generalization.
@@
In addition to dataset scale, there is a need for novel neural architectures and training strategies tailored to model audio structures more effectively. For instance, LAION-CLAP [16] investigates different choices of audio/text encoders and achieves the best performance using HTSAT [19] for audio and RoBERTa [20] for text. In addition, LAION-CLAP proposes feature fusion for audio inputs with variable-length. In concurrent work, Fast Language-Audio Pretraining (FLAP) [21], inspired by Fast Language-Image Pretraining (FLIP) [22], proposes masking and removing a significant portion of spectrogram patches during contrastive training. FLAP also incorporates a reconstruction loss during the contrastive training on these masked spectrogram patches, although the improvements are modest.
@@
In this paper, we investigate several strategies to improve the CLAP models, informed by the aforementioned challenges. For dataset creation, we collect a large-scale audio-text dataset and expand and refine its text descriptions. For audio recordings with weak or no labels, we utilize an automatic audio captioning model to obtain synthetic captions. For audio paired with noisy descriptions, we use LLMs to generate several cleaned captions for each audio clip. The result of these efforts leads to a collection of over 3.9 million text-audio pairs, with over 13,000 hours of audio.
@@
For our neural architecture and training strategy, we propose to use a two-stage approach. The first stage focuses on training spectrogram-based audio encoder using a masked autoencoder (MAE) objective [23, 24]. In this stage, MAE learns representations through masking random patches from the input spectrogram and then reconstructing these missing patches. We anticipate that the MAE pre-pretraining will provide a significantly better initialization for the following contrastive training. A classification objective, on the other hand, may allow the model to discard important information that is not necessary for classification. In the second stage, we use the audio encoder from the first stage to train our audio-text model on collected synthetic text-audio pairs, employing dual contrastive and captioning objectives. The integration of the auxillary captioning objective, inspired by CoCa [25] and BLIP [26], provides stronger supervision, encouraging the audio encoder to capture fine-grained patterns that closely match text descriptions. Training a captioner decoder also facilitates text generation for open-ended audio understanding tasks, expanding our model’s application scope.
@@
In the evaluation phase, incorporating a diverse range of evaluation tasks is crucial to comprehensively measuring model capability and preventing overfitting to common test sets, as suggested by Recht et al. [27]. Typically, audio-text representation learning is assessed through zero-shot audio classification and audio-text retrieval. To provide a more comprehensive benchmark, we additionally evaluate on audio question answering (AQA) [28]. Compared to text-audio retrieval tasks or classification tasks, AQA aims at capturing fine-grained details from the sound events, such as temporal dynamics between audio and text [29]. To evaluate the effectiveness of our audio encoder, we test on Holistic Evaluation of Audio Representations (HEAR) [30]. HEAR offers broad evaluation tasks that test the general-purpose audio representation through audio classification and sound event detection. Lastly, to assess the performance of our captioning decoder, we evaluate on automatic captioning tasks for open-ended generation.
@@
In summary, the contribution of our work includes: (1) We curate a large-scale refined audio-text dataset with LLM processing and audio captioning. (2) We propose a two-stage training approach for contrastive models: we first pre-pretrain an audio encoder with an MAE objective. In the second stage, we train a constrative model, initializing the audio encoder from the first stage, and include an auxiliary captioning objective to enhance the model’s understanding of audio-text relationships. (3) We benchmark our model on a variety of audio understanding tasks. Particularly, our model achieves state-of-the-art or equivalent performance on audio-text retrieval tasks. We have also conducted comprehensive ablation studies to demonstrate the impact of our different contributions. We open source the inference and evaluation codebase along with our pretrained model [^1] https://github.com/gzhu06/Cacophony.
@@ II Dataset Collection
Figure 2: Overview of our proposed dataset creation pipeline. Left: We process text descriptions based on the dataset quality. It aims to automatically clean and generate audio captions while maintaining consistency between the audio content and the textual descriptions. Middle: For datasets with raw, noisy descriptions, we utilize cascaded fine-tuning with large language models, GPT-3 and T5-large, to clean the information that is irrelevant to the sound. Right: The detailed process of description cleaning using GPT-3 based on few-shot prompting is outlined as follows: Freesound raw inputs are highlighted in blue, sample inputs provided by a human annotator are marked in red, and the output text generated by GPT-3 is shown in bold red.
We categorize the publicly available audio datasets into three types based on label granularity: clean-labeled, noisy-labeled, and weakly-labeled or unlabeled. The number of samples of each category is shown in Fig. 1. Clean-labeled datasets typically contain high-quality, manually curated text descriptions, but are often smaller in size. Noisy-labeled datasets include text descriptions that provide relevant sound information but also tend to include extraneous details. Weakly-labeled and unlabeled datasets vary from having incomplete to no textual information about the recordings. As illustrated in Fig. 1, it can be observed that the total size of all combined clean-labeled datasets is small compared to the noisy datasets, and the weakly-labeled / unlabeled category dwarfs both the clean-labeled and noisy-labeled categories. To make use of the audio data with either noisy or missing labels, we leverage publicly available tools to synthesize audio descriptions, shown in the left of Fig. 2. In the remainder of this section, we propose strategies to effectively handle each specific case.
@@ II-A Clean-labeled datasets
The ‘clean-labeled’ dataset split, which we refer to as ‘OpenSFX’, comprises AudioCaps-train, Clotho-development, Epidemic Sound, SoundDescs, Free To Use Sounds, Sonniss Game Effects, AudioStock, and MACS. The most widely used datasets in this category are AudioCaps [14] and Clotho [15]. AudioCaps contains approximately 50,000 audio clips sourced from AudioSet [17], and are annotated by humans. Clotho includes around 6,000 audio clips from Freesound222www.freesound.org, with each clip featuring five human-generated captions. In addition to the commonly referenced datasets, there are others like the SoundDescs dataset [31], WavText5K[32], and Epidemic Sound333www.epidemicsound.com. The “writing styles” of the captions differ across audio-text datasets. For instance, Audiostock has annotations of relatively higher quality, similar to those in AudioCaps, such as “The sound of closing the zipper on the makeup pouch.” Meanwhile, the captions in SoundDescs are full of entities, for example, “Wolseley Sixteen-sixty Car 1967 (Exterior) - Exterior, approach and stop, tickover, switch off,” and Epidemic Sound consists a sequence of tags: ”the sounds of pop, hollow, airy, smacking impacts, sharp, small, punchy, and soft hits.” While the annotation formats vary across these datasets, their raw descriptions still offer detailed content information about the audio clips. In our case, aiming to incorporate as much data as possible, we opt not to implement text-cleaning steps for processing these captions.
@@ II-B Noisy-labeled datasets
We also include a larger-scale audio data source, Freesound, which contains noisy text descriptions. Freesound is a collaborative online platform dedicated to sharing sounds, hosting over 500k audio clips uploaded by users, and has been used in studies such as WavCaps [18] and LAION-CLAP [16]. These clips, varying in duration, cover a wide range of audio content including music, environmental sounds, synthesized effects, and various noises. We exclude clips that are under one second or exceed five minutes, as shorter clips typically offer limited meaningful content and require excessive padding for training, and longer clips are often redundant. Freesound prompts users to give brief descriptions for each audio clip they upload. However, these descriptions are often inaccurate and sometimes include named entities such as people’s names, locations, and details about recording equipment, as demonstrated by the example description provided on the right shown in Fig. 2.
We propose to use LLMs to transform raw Freesound descriptions into usable captions by automatically removing sound-irrelevant information. Although ChatGPT is well-suited for this task, the cost of generating captions for each Freesound sample would be prohibitively expensive. Given our resources, it would be more feasible to use a T5 model  [33], but we find that available pretrained T5 models struggle with this task. We thus use GPT-3.5  [^4] Model name: gpt-3.5-turbo-0301 to construct a small dataset of raw-clean pairs, which will then be used to fine-tune T5-XXL [33] specifically for caption cleaning. This fine-tuned T5 model is subsequently used to generate clean captions for the entire Freesound dataset, as illustrated in the middle section of Fig. 2.
More specifically, to generate the synthetic dataset used for fine-tuning T5, we few-shot prompt GPT-3.5 to generate a clean caption given a noisy one with 10 human annotated pairs of noisy/clean captions in our prompt. This is shown in the right section of Fig. 2. We take a random sample of 7,000 (out of the total 500,000) Freesound noisy text captions and use GPT-3.5 to generate corresponding clean captions.
After fine-tuning the T5-XXL model with these pairs, we generate text descriptions for each audio sample in Freesound. We generate five descriptions per audio sample, which serves as a form of data augmentation. Tab. I presents examples of the Freesound raw descriptions and LLM-processed captions. Our fine-tuned T5 model is capable of automatically translating Spanish to English, eliminating redundant or irrelevant details, and summarizing long sentences into one-sentence high-level audio captions. However, the model inaccurately represents the audio recordings in a small portion of the captions. For instance, some captions mistakenly identify ‘bike’ as ‘motorbike’ or fail to entirely remove entities, as exemplified at the bottom of Tab. I.
@@
II-C Weakly-labeled/unlabeled datasets
Lastly, the datasets with largest scale are those that are weakly-labeled or unlabeled, including AudioSet [17] and ACAV100M [34]. Both AudioSet and ACAV100M are comprised of audio clips extracted from YouTube videos, offering a vast range of audio data. AudioSet is a weakly-labeled dataset with 527 predefined sound classes of 2 million 10-second recordings, where each audio clip is marked only by the presence of sound event tags, without detailed descriptions. ACAV100M consists of 100 million videos that exhibit high audio-visual correspondence, yet it lacks of any form of labeling. The ACAV dataset also contains a variety of scales, ranging from 20K to 100M. In this study, we concentrate on ACAV2M, leaving the exploration of larger scales for future work. To leverage these datasets and to accurately represent audio content, we employ an off-the-shelf audio captioning model HTSAT-BART pretrained on WavCaps and fine-tuned on AudioCaps  [^5] https://github.com/XinhaoMei/WavCaps. Specifically, we choose this model because it achieves state-of-the-art performance on audio captioning tasks and is trained on a diverse range of data. Employing the chosen captioning model for synthetic caption generation differs from previous work such as BLAT [35], whose captioner is initially trained on the smaller-scale AudioCaps dataset.

### v3

@@ Abstract
Despite recent advancements, audio-text models still lag behind their image-text counterparts in scale and performance. In this paper, we propose to improve both the data scale and the training procedure of audio-text contrastive models. Specifically, we craft a large-scale audio-text dataset containing 13,000 hours of text-labeled audio, using pretrained language models to process noisy text descriptions and automatic captioning to obtain text descriptions for unlabeled audio samples. We first train on audio-only data with a masked autoencoder (MAE) objective, which allows us to benefit from the scalability of unlabeled audio datasets. We then train a contrastive model with an auxiliary captioning objective with the audio encoder initialized from the MAE model. Our final model, which we name Cacophony, achieves state-of-the-art performance on audio-text retrieval tasks, and exhibits competitive results on the HEAR benchmark and other downstream tasks such as zero-shot classification.
@@ I Introduction
Machine audition [1] involves developing algorithms and systems for machines to analyze and understand sound, covering tasks such as audio tagging, acoustic scene classification, music classification, and sound event detection. In recent years, there has been a general shift away from individual audio pattern recognition tasks and toward general-purpose audio representations pretrained on large-scale audio datasets. Pretrained Audio Neural Networks (PANNs)  [2] have played a significant role in this shift by demonstrating their versatility across various tasks and outperforming many advanced systems via fine-tuning.
@@
Modern approaches aim for robust performance in general-purpose audio understanding tasks without the need for task-specific fine-tuning, which offers more flexibility. These methods approach general audio understanding by linking text and audio modalities, referred to as audio-text models. One approach to linking text and audio is through generating response text given a combination of an audio prompt and a text prompt [3, 4, 5]. For instance, Pengi [5] converts audio classification, retrieval, captioning, and audio question answering into a text generation task using audio and task-specific text prompts. Similarly, Qwen-Audio [6] addresses a variety of audio tasks through text generation but distinguishes itself by using text prompts consisting of hierarchical tag sequences inspired by Whisper [7]. Both Pengi and Qwen-Audio support multiple closed-ended and open-ended audio tasks without the need for additional fine-tuning or task-specific extensions of the architecture. Another method of linking text and audio is contrastive learning. Pretrained contrastive models can also be applied directly to various downstream tasks without fine-tuning. For instance, contrastive models can be used for retrieval and classification by assigning a score that identifies the most probable text (or class label) from a predefined set of choices for a given audio input. Moreover, the learned audio-text representations from contrastive models offer the flexibility to use one modality during training and the other at inference, which can be applied in text-to-audio generation [8, 9] and language-guided source separation [10, 11].
In this paper, we focus on improving contrastive audio-text models for general sounds by addressing two critical limitations in existing research: insufficient dataset scale and the vanilla contrastive training techniques. To tackle the dataset scale issue, we first examine the publicly available audio datasets. These can be categorized into three types based on label granularity: clean-labeled, noisy-labeled, and weakly-labeled or unlabeled, as illustrated in Fig. 1. Clean-labeled datasets, while high-quality, are limited in size. Noisy-labeled datasets offer more samples but include extraneous details. The largest category by far is weakly-labeled or unlabeled data, which provides little to no textual information. This scarcity of high-quality labeled data presents a significant challenge in audio-text contrastive learning, especially when compared to recent advancements in image-text models. For context, while the largest public audio-text datasets contain less than 100,000 pairs [12, 13], image-text models like CLIP [14] and SigLip [15] utilize 400 million and 3.6 billion pairs respectively.
@@
Previous works have attempted to address this data scarcity issue by collecting data from various sources and applying natural language processing techniques to clean or filter noisy captions. For instance, Huang et al. [16] collected approximately 44 million 30-second music clips, applying a pretrained classifier and rule-based filtering to clean associated metadata. This process intensively reduced the dataset to 2 million music-text pairs. Notably, they found that models trained on large scale unfiltered audio-text data performed comparably to those trained on filtered data in music tagging and retrieval tasks, suggesting that data quantity might be as crucial as quality in this domain. Wu et al. [17], in collaboration with Large-scale Artificial Intelligence Open Network (LAION), curate the LAION-Audio dataset with 630K audio-text pairs together with 2 million Audioset [18] clips recaptioned with keyword-to-caption (K2C) augmentation to train LAION-CLAP. The K2C method uses a pretrained language model to generate captions from tags. However, the captions produced via K2C are restricted to the objects defined by the tags, offering limited descriptive details. K2C also risks making incorrect assumptions or introducing biases highlighted in [17]. Mei et al. [19] propose a multi-stage data filtering pipeline and utilize ChatGPT for cleaning text descriptions. Their contrastive language-audio pretraining (CLAP) models, trained on their WavCaps dataset, demonstrate superior performance in audio-text retrieval tasks compared to LAION-CLAP, despite a smaller-scale dataset. Such a text filtering pipeline reduces the amount of audio data, which could potentially result in reduced generalization.
@@
In addition to dataset scale, there is a need for novel neural architectures and training strategies tailored to model audio structures more effectively. For instance, LAION-CLAP [17] investigates different choices of audio/text encoders, demonstrating superior performance with the hierarchical token semantic audio transformer (HTSAT)[20] for audio encoding and the Robustly optimized BERT111acronym for ‘Bidirectional Encoder Representations from Transformers’ approach (RoBERTa)[21] for text encoding. In addition, LAION-CLAP proposes feature fusion for audio inputs with variable-length. In concurrent work, fast language-audio pretraining (FLAP) [22], inspired by fast language-image pretraining (FLIP) [23], proposes masking and removing a significant portion of spectrogram patches. FLAP also incorporates a reconstruction loss during the contrastive training on these masked spectrogram patches, although the improvements are modest.
@@
In this paper, we investigate several strategies to improve the audio-text models, informed by the aforementioned challenges. For dataset creation, we collect a large-scale audio-text dataset and expand and refine its text descriptions. For audio recordings with weak or no labels, we utilize an automatic audio captioning model to obtain synthetic captions. For audio paired with noisy descriptions, we use large language models (LLMs) to generate several cleaned captions for each audio clip. These efforts lead to a collection of over 3.9 million audio-text pairs, with over 13,000 hours of audio.
@@
For our neural architecture and training strategy, we propose to use a two-stage approach. The first stage focuses on training spectrogram-based audio encoder using a masked autoencoder (MAE) objective [24, 25], which learns representations through masking random patches from the input spectrogram and then reconstructing these masked patches. We anticipate that the MAE training will provide a better initialization for the following contrastive training. An audio classification objective, in contrast, may encourage the model to discard information unnecessary for classification but important for contrastive training. In the second stage, we use the audio encoder from the first stage to train our audio-text model on collected synthetic audio-text pairs, employing dual contrastive and captioning objectives. The integration of the auxillary captioning objective, inspired by contrastive captioner (CoCa) [26] and bootstrapping language-image pre-training (BLIP) [27], provides stronger supervision, encouraging the audio encoder to capture fine-grained patterns that closely match text descriptions. Training a captioner decoder also facilitates text generation for open-ended audio understanding tasks, expanding our model’s application scope.
@@
In the evaluation phase, incorporating a diverse range of evaluation tasks is crucial to comprehensively measuring model capability and preventing overfitting to common test sets, as suggested by Recht et al. [28]. Typically, audio-text representation learning is assessed through zero-shot audio classification and audio-text retrieval. To provide a more comprehensive benchmark, we additionally evaluate on audio question answering (AQA) [29]. To evaluate the effectiveness of our audio encoder, we test on Holistic Evaluation of Audio Representations (HEAR) [30]. HEAR offers broad evaluation tasks that test the general-purpose audio representation through audio classification and sound event detection. Lastly, to assess the performance of our captioning decoder, we evaluate on automatic captioning tasks for open-ended generation.
@@
In summary, the contribution of our work includes: (1) We curate a large-scale refined audio-text dataset with LLM processing and audio captioning. (2) We propose a two-stage training approach for contrastive models: we first train an audio encoder with an MAE objective. In the second stage, we train a constrative model, initializing the audio encoder from the first stage, and include an auxiliary captioning objective to enhance the model’s understanding of audio-text relationships. (3) We benchmark our model on a variety of audio understanding tasks. Particularly, our model achieves state-of-the-art or comparable performance on audio-text retrieval tasks. We have also conducted comprehensive ablation studies to demonstrate the impact of our different contributions. We open source the inference and evaluation codebase along with our pretrained model [^2] https://github.com/gzhu06/Cacophony.
@@ II Dataset Collection
Figure 2: Overview of our proposed dataset creation pipeline. Left: We process text descriptions based on the dataset quality. It aims to automatically clean and generate audio captions while maintaining consistency between the audio content and the textual descriptions. Middle: For datasets with raw, noisy descriptions, we use large language models, GPT-3 and T5-XXL, to clean the information that is irrelevant to the sound. Right: The detailed process of description cleaning using GPT-3 based on few-shot prompting is outlined as follows: Freesound raw inputs are highlighted in blue, sample inputs provided by a human annotator are marked in red, and the output text generated by GPT-3 is shown in bold red.
To make use of the audio data with either noisy or missing labels, we leverage publicly available tools to synthesize audio descriptions, shown in the left of Fig. 2. In the following, we propose strategies to effectively handle each case.
@@ II-A Clean-labeled datasets
The “clean-labeled” dataset split, which we refer to as “OpenSFX”, comprises AudioCaps-train, Clotho-development, Epidemic Sound, SoundDescs, Free To Use Sounds, Sonniss Game Effects, AudioStock, and MACS. The most widely used datasets in this category are AudioCaps [12] and Clotho [13]. AudioCaps contains approximately 50,000 audio clips sourced from AudioSet [18] and annotated by humans. Clotho includes around 6,000 audio clips from Freesound333www.freesound.org, with each clip featuring five human-generated captions. In addition to the commonly referenced datasets, there are others like the SoundDescs dataset [31], WavText5K[32], and Epidemic Sound444www.epidemicsound.com. The “writing styles” of the captions differ across audio-text datasets. Nevertheless, their raw descriptions all offer detailed content information about the audio clips. In our case, aiming to incorporate as much data as possible, we opt not to implement text-cleaning steps for processing these captions.
@@ II-B Noisy-labeled datasets
We also include a larger-scale audio data source, Freesound, which contains noisy text descriptions. Freesound is a collaborative online platform dedicated to sharing sounds, hosting over 500k audio clips uploaded by users, and has been used in studies such as WavCaps [19] and LAION-CLAP [17]. These clips, varying in duration, cover a wide range of audio content including music, environmental sounds, synthesized effects, and various noises. We exclude clips that are under one second or exceed five minutes, as shorter clips typically offer limited meaningful content and require excessive padding for training, and longer clips are often redundant. While Freesound prompts data uploaders give a brief description for each audio clip, these descriptions are often inaccurate and sometimes include named entities such as people’s names, locations, and details about recording equipment.
We propose to use LLMs to transform raw Freesound descriptions into usable captions by automatically removing sound-irrelevant information. Although ChatGPT is well-suited for this task, the cost of generating captions for each Freesound sample would be prohibitively expensive. Given our resources, it would be more feasible to use a T5 model  [33], but we find that available pretrained T5 models struggle with this task. We thus use GPT-3 to construct a small dataset of raw-clean pairs [^5] Model name: gpt-3.5-turbo-0301, which will then be used to fine-tune T5-XXL [33] specifically for caption cleaning. This fine-tuned T5 model is subsequently used to transform the raw Freesound descriptions to clean captions for the entire Freesound dataset, as illustrated in the middle section of Fig. 2.
More specifically, to generate the synthetic dataset used for fine-tuning T5, we few-shot prompt GPT-3 to generate a clean caption given a noisy one with 10 human annotated pairs of noisy/clean captions in our prompt. This is shown in the right section of Fig. 2. We take a random sample of 7,000 (out of the total 500,000) Freesound noisy text captions and use GPT-3 to generate corresponding clean captions.
After fine-tuning the T5-XXL model with these pairs, we cleaned the text description for each audio sample in Freesound and generate five clean version; this servers as a form of data augmentation.
Table I presents examples of the Freesound raw descriptions and LLM-processed captions. Our fine-tuned T5 model is capable of automatically translating Spanish to English, eliminating redundant or irrelevant details, and summarizing long sentences into one-sentence high-level audio captions. However, the model inaccurately represents the audio recordings in a small portion of the captions. For instance, some captions mistakenly identify ‘bike’ as ‘motorbike’ or fail to entirely remove entities, as exemplified at the bottom of Table I.
@@
II-C Weakly-labeled or unlabeled datasets
Lastly, the datasets with the largest scale are those that are weakly-labeled or unlabeled, including AudioSet [18] and ACAV100M [34]. Both AudioSet and ACAV100M are comprised of audio clips extracted from YouTube videos, offering a vast range of audio data. AudioSet is a weakly-labeled dataset with 527 predefined sound classes of 2 million 10-second recordings, where each audio clip is marked only by the presence of sound event tags, without detailed descriptions. ACAV100M consists of 100 million videos that exhibit high audio-visual correspondence, yet it lacks of any form of labeling. The ACAV dataset also contains a variety of scales, ranging from 20K to 100M. In this study, we concentrate on ACAV2M, leaving the exploration of larger scales for future work. To leverage these datasets and to accurately represent audio content, we employ an off-the-shelf audio captioning model HTSAT-BART pretrained on WavCaps and fine-tuned on AudioCaps  [^6] https://github.com/XinhaoMei/WavCaps. Specifically, we choose this model because it achieves state-of-the-art performance on audio captioning tasks and is trained on a diverse range of data. Employing the chosen captioning model for synthetic caption generation differs from previous work such as BLAT [35], whose captioner is initially trained on the smaller-scale AudioCaps dataset.